# State-Space Modeling in Long Sequence Processing: A Survey on Recurrence in the Transformer Era

### v1

@@ 7.5 Time Series Benchmarks
Time series consists of sequential samples paired with time-stamps, generally (but not necessarily) sampled at equally spaced points in time. Such data is often characterized by very long range recurring patterns, that can be exploited to extract meaningful statistics. Several problems can be faced with these data, such as ($i$) the classification of a time series based on its characteristics; ($ii$) regression/prediction analysis to test relationships between one or more different time series; ($iii$) time series forecasting, searching for a model capable of predicting future values of the series based on previously observed ones.
@@
Regression/Prediction.   The BDIMC healthcare datasets is composed by $7,949$ sequences, and it is aimed at predicting three vital signs of a patient. It is part of the TSR archive [281], presenting clinical data from the Beth Israel Deaconess Medical Center. PPG and ECG signals were sampled with a frequency of $125$Hz for $8$ minutes each. The resulting two-dimensional sequences have a length of $L=4000$. The goal is to predict a patient’s respiratory rate (RR), heart rate (HR), and oxygen saturation (SpO2), based on PPG and ECG signals. RMSE is exploited to measure performances [249, 90].
@@ 7.7 Long Range Arena
The previous sections highlighted the absence of an established common set of benchmarks to evaluate sequential model that handles long range tasks. Indeed, when comparing the experimental setup of the previously described papers, it can be observed that there is not a consensus on a common testbed. Models are evaluated on heterogeneous tasks and datasets (even when using the same dataset, different works often preprocess it in a different manner), making the comparison of their performances very difficult. Moreover, many of the benchmarks are not directly targeting long-range modeling ability. Finally, recent works fall short in decoupling the effect of large scale pretraining from the inductive bias emerging from a certain data distribution. In an effort to standardize long-range models evaluation, Tay et al. [282] proposed the Long Range Arena (LRA) benchmark suite. LRA includes synthetic and real-world tasks to assess the ability of sequential architectures to reason in such scenarios, considering different types of data and conditions. It focuses on understanding how well architectures can handle long sequences having certain hierarchical or spatial structures, and aims to compare their performance across various settings. In details, it consists of $6$ tasks: ($i$) Long ListOps, ($ii$) Text ($iii$) Retrieval ($iv$) Image ($v$) Pathfinder and its extreme long version, ($vi$) Path-X. Sequences constituting these task have a length $L\in[1k,16k]$, encompassing modalities and objectives that require similarity, structural, and visuospatial reasoning. The LRA suite is becoming increasingly popular in the scientific literature and has been exploited by several recent works mentioned in the context of this survey [88, 94, 87, 271, 100, 171, 235, 233, 170, 230, 224, 320, 213, 95, 159].
@@
Input symbols are encoded via one-hot vectors, having $17$ unique possible configurations (opening brackets and operators are grouped into a single token). Long ListOps tests the ability to reason hierarchically while handling long contexts with varying length, up to $L=2$k, with a reserved end-of-sequence token. The final goal is ten-way classification (the integer result of the expression). There are $96$k training sequences, $2$k validation sequences, and $2$k test sequences.
@@
Text.   Text is a byte-level text classification task, based upon the IMDb binary sentiment classification previously described (i.e., classify whether a movie review is positive/negative). Text sequences are processed at byte/character-level (characters are encoded as one-hot vectors, with $129$ unique values possible), with a fixed max length of $L=4$k, and truncated or padded when necessary. The evaluation metric is the accuracy. Compared to ListOps, the model needs to reason with compositional unsegmented real-world data, with less defined boundaries, that must be learnt from data. Differently from character-level language modeling, where the local context is sufficient to infer the next character, here the model is required to grasp the overall sentiment of the processed text, hence it needs to compose characters into words and then higher-level concepts. There are $25$k training examples and $25$k test examples, while no validation set is provided.
@@
Retrival.   Retrieval can be used to assess the model capability of encoding long sequences into representations suitable for similarity-based matching. Couples of texts are compressed and concatenated, then provided to a linear classifier. The text is sampled from the ACL Anthology Network [236] dataset. Given two textual citations, where characters are encoded as a sequence of integer tokens (one-hot vector with $97$ unique values), the task is to classify whether the two citations are equivalent. Byte/character level processing is performed, resulting in sequence length of $L=4$k for each document (i.e., total text length $8$k for this task). The performances of this binary classification task are measured with accuracy. There are $147,086$ training pairs, $18,090$ validation pairs, and $17,437$ test pairs.
@@
Image.   The Image task corresponds to the already introduced sCIFAR, an image classification task on sequences of pixels, described in Section 7.2. A $32\times 32$ pixels image is flattened in raster scan into a sequence of $1024$ pixels. The goal is to classify ($10$-way classification) the image by processing the 1D pixel sequence. Models are required to capture 2D spatial hierarchical structure between pixels even if they are processed sequentially. Images belong to the CIFAR-10 [147] dataset, and are processed in a single gray-scale channel (each pixel represented with 8-bit pixel intensity, hence vocabulary size of $256$). [^30] Some recent works process the coloured images instead [213]. There are $45$k training examples, $5$k validation examples, and $10$k test examples.
@@
A $32\times 32$ grayscale image contains two points (a start and an end point, represented as small circles—see Figure 19). There are dashed lines over the image. The task is to make a binary decision whether the two points are connected by a path (Figure 19-left) or not (Figure 19-right). Models process the image pixel sequence (raster scan), thus $L=1024$. There are $160$k training examples, $20$k validation examples, and $20$k test examples.
@@
Path-X.   Finally, Path-X is an harder version of Pathfinder composed by $128\times 128$ images, resulting in sequences of $L=16$k.
@@
Benchmarks Sec. 3 of this paper Sec. 4 of this paper Sec. 5 of this paper Sec. 6 of this paper
Synthetic Adding Problem [105] [150, 5] - [157, 89, 248, 250] [130]
Copy Task [105] [5, 191, 10, 30, 58] [86, 320, 57, 85] [89, 140] [191, 117, 217]
@@
Associative Recall [309, 30, 223] [55, 85, 227] [15] [162]
Computer Vision sMNIST [293, 86, 90, 88, 271] [89, 140, 248, 164, 68, 250] [130, 211]
psMNIST [248, 249, 68, 250] [130, 217]
@@
Natural Language WikiText103 [192] [224, 118, 259, 230, 111, 51, 30, 158, 133, 180, 233, 304, 135, 235] [88, 171, 159, 227] [89, 15] [191]
PTB[181] [51] [140, 250, 157, 25, 131] [130]
@@
Time Series Eigenworms [27] [116, 249, 165]
@@ 8 Discussion and Future Directions
In this Section, we analyze some open problems and issues that have been recently pointed out in the works described in this survey, as long as highlighting possible future avenues for research.
8.1 The Attention-Recurrence Expressivity Gap
@@
In the context of this survey we analyzed several architectural solutions based on recurrence, leveraging its stateful nature. Indeed, recent models exhibit constant-time inference and linear-time training (e.g, in SSMs [86, 90, 271], see Section 4) providing benefits when compared to stateless methods whose complexity is quadratic in the sequence length, such as vanilla Transformers. However, despite the computational advantages of recurrence-based models, performances are not always on par with respect to vanilla attention when considering specific tasks, also in synthetic settings [55, 124]. Indeed, the capability of recurrent models to effectively solve a target task is strongly conditioned upon how efficiently the processed sequence is compressed into a finite-size state [85], which acts as an information bottleneck.
@@
The role of gating mechanisms turns out to be fundamental, given their influence on the model ability to control, in a context/input dependant manner, how information propagates or interact along the temporal dimension. Indeed, the relevance of an appropriate gating mechanism can be appreciated when considered that such a mechanism is present in many high-impact works described in this survey [126, 309, 319, 259, 118, 222, 277, 135], going well beyond the original case of LSTMs. Indeed, the Mamba model [85] shows how a selection mechanism (i.e., a term which refers to the mechanistic action of a model to select or ignore inputs and facilitate data interaction along the sequence length, inspired by the original gating in [105]) improves the capability of State-Space Models to solve tasks such as Selective copy or Induction Heads (see Section 7). In these task, content-aware reasoning is fundamental to be able to memorize relevant tokens and filter out the irrelevant ones. While the Linear Time Invariant (LTI) nature of previous SSMs (LSSL, S4 [90, 88]) allows to achieve efficient parallelizable solutions, their dynamics do not easily allow to select the correct information from the current context, making it harder to encode the information stored into the hidden state. Conversely, an appropriate selection mechanism introduces context-awareness, allowing the model to focus on or to filter out information. This element is also interesting in the light of recent studies (see Section 5) suggesting that certain RNNs (equipped with linear recurrent layers interconnected by feed-forward paths with multiplicative gating) might be implementing attention/selection “under the hood”, when trained to solve simple in-context learning tasks [319].
@@
However, despite such advantages brought by gating/selection mechanisms, performances of recurrence-based solutions are still not capable to fully close the gap with attention-based models. Several recent works investigated such a recurrence expressivity gap [55, 124] with respect to attention. Indeed, inference in Transformers leverages the entire sequence processed so far, which, in autoregressive models, is stored in the KV-cache (see Section 3), to produce the final output. By definition, the memory usage of this approach increases over time, which becomes extremely expensive. At the cost of inefficiency, the attention mechanism posses properties that standard stateful models are not able to embrace: explicitly recalling earlier tokens in the sequence (that influences their ability to copy from the input sequence, possibly by multiplying the attention scores with $V$) and comparing tokens across the sequence (which is achievable thanks to the quadratic attention matrix $QK^{\prime}$)). Jelassi et al. [124] exploit the copy task (see Section 7) to theoretically and empirically prove this property. In short, they show that from the theoretical point of view a simple Transformer model could be capable (when empowered by proper positional encodings and under some other mild assumptions) to copy strings of length that is exponential in the number of heads of the Transformer. This result relies on the ability of the Transformer to implement a mechanism of “storage” and retrieval of sequences of $n$ tokens ($n$-grams). Noticeably, the number of parameters depends only logarithmically on the input sequence length. The authors show that, conversely, any state-space model fails to solve the copy task, unless its latent state grows linearly with the sequence length. That is, SSMs cannot accurately copy strings with more bits than the size of the latent state.
@@
Such theoretical findings have been confirmed empirically. Indeed, Transformers are both much more efficient at learning to copy (i.e., they need $\approx 10^{4}$ training examples to learn the task, whereas Mamba [85] needs more than $\approx 10^{6}$ samples) and also generalize better to longer inputs (when trained on strings $L=50$, they generalize to sequences with $L>1000$, whilst Mamba fails to generalize out of the trained context). This finding confirm the theoretical results, i.e., to be able to copy the entire input sequence, stateful models need to fully store it in its state space, which requires the memory to grow linearly with the sequence length. These results hold even when considering pre-trained models, both in copy tasks and retrieval benchmarks (e.g., Phonebook Lookup, see Section 7).

### v2

@@ 7.5 Time Series Benchmarks
Time series consists of sequential samples paired with time-stamps, generally (but not necessarily) sampled at equally spaced points in time. Such data is often characterized by very long range recurring patterns, that can be exploited to extract meaningful statistics. Several problems can be faced with these data, such as ($i$) the classification of a time series based on its characteristics; ($ii$) regression/prediction analysis to test relationships between one or more different time series; (iiiiii) time series forecasting, searching for a model capable of predicting future values of the series based on previously observed ones.
@@
Regression/Prediction.   The BDIMC healthcare datasets is composed by $7,949$ sequences, and it is aimed at predicting three vital signs of a patient. It is part of the TSR archive [293], presenting clinical data from the Beth Israel Deaconess Medical Center. PPG and ECG signals were sampled with a frequency of 125125Hz for $8$ minutes each. The resulting two-dimensional sequences have a length of $L=4000$. The goal is to predict a patient’s respiratory rate (RR), heart rate (HR), and oxygen saturation (SpO2), based on PPG and ECG signals. RMSE is exploited to measure performances [186, 40].
@@ 7.7 Long Range Arena
The previous sections highlighted the absence of an established common set of benchmarks to evaluate sequential models that handle long-range tasks. Indeed, when comparing the experimental setup of the previously described papers, it can be observed that there is not a consensus on a common testbed. Models are evaluated on heterogeneous tasks and datasets ; even when using the same dataset, different works often preprocess it in a different manner, making the comparison of their performances very difficult. Moreover, many of the benchmarks do not directly target long-range modeling ability. Finally, recent works fall short in decoupling the effect of large-scale pretraining from the inductive bias arising from a certain data distribution.
In an effort to standardize long-range models evaluation, Tay et al. [35] proposed the Long Range Arena (LRA) benchmark suite. LRA includes synthetic and real-world tasks to assess the ability of sequential architectures to reason in such scenarios, considering different types of data and conditions. It focuses on understanding how well architectures can handle long sequences having certain hierarchical or spatial structures, and aims to compare their performance across various settings. In details, it consists of $6$ tasks: ($i$) Long ListOps, ($ii$) Text (iiiiii) Retrieval ($iv$) Image ($v$) Pathfinder and its extreme long version, ($vi$) Path-X. Sequences constituting these task have a length $L\in[1k,16k]$, encompassing modalities and objectives that require similarity, structural, and visuospatial reasoning. The LRA suite is becoming increasingly popular in the scientific literature and has been exploited by several recent works mentioned in the context of this survey [25, 149, 150, 42, 299, 135, 131, 102, 101, 103, 84, 212, 43, 41, 154].
@@
Input symbols are encoded via one-hot vectors, having $17$ unique possible configurations (opening brackets and operators are grouped into a single token). Long ListOps tests the ability to reason hierarchically while handling long contexts with varying length, up to L=2L=2k, with a reserved end-of-sequence token. The final goal is ten-way classification (the integer result of the expression). There are 9696k training sequences, 22k validation sequences, and 22k test sequences.
@@
Text.   Text is a byte-level text classification task, based upon the IMDb binary sentiment classification previously described (i.e., classify whether a movie review is positive/negative). Text sequences are processed at byte/character-level (characters are encoded as one-hot vectors, with $129$ unique values possible), with a fixed max length of L=4L=4k, and truncated or padded when necessary. The evaluation metric is the accuracy. Compared to ListOps, the model needs to reason with compositional unsegmented real-world data, with less defined boundaries, that must be learnt from data. Differently from character-level language modeling, where the local context is sufficient to infer the next character, here the model is required to grasp the overall sentiment of the processed text, hence it needs to compose characters into words and then higher-level concepts. There are 2525k training examples and 2525k test examples, while no validation set is provided.
@@
Retrieval.   Retrieval can be used to assess the model capability of encoding long sequences into representations suitable for similarity-based matching. Couples of texts are compressed and concatenated, then provided to a linear classifier. The text is sampled from the ACL Anthology Network [301] dataset. Given two textual citations, where characters are encoded as a sequence of integer tokens (one-hot vector with $97$ unique values), the task is to classify whether the two citations are equivalent. Byte/character level processing is performed, resulting in sequence length of L=4L=4k for each document (i.e., total text length 88k for this task). The performances of this binary classification task are measured with accuracy. There are $147,086$ training pairs, $18,090$ validation pairs, and $17,437$ test pairs.
@@
Image.   The Image task corresponds to the already introduced sCIFAR, an image classification task on sequences of pixels, described in Section 7.2. A $32\times 32$ pixels image is flattened in raster scan into a sequence of $1024$ pixels. The goal is to classify ($10$-way classification) the image by processing the 1D pixel sequence. Models are required to capture 2D spatial hierarchical structure between pixels even if they are processed sequentially. Images belong to the CIFAR-10 [246] dataset, and are processed in a single gray-scale channel (each pixel represented with 8-bit pixel intensity, hence vocabulary size of 256256). [^31] Some recent works process the coloured images instead [43].. There are 4545k training examples, 55k validation examples, and 1010k test examples.
@@
A $32\times 32$ grayscale image contains two points (a start and an end point, represented as small circles—see Figure 19). There are dashed lines over the image. The task is to make a binary decision whether the two points are connected by a path (Figure 19-left) or not (Figure 19-right). Models process the image pixel sequence (raster scan), thus $L=1024$. There are 160160k training examples, 2020k validation examples, and 2020k test examples.
@@
Path-X.   Finally, Path-X is an harder version of Pathfinder composed by $128\times 128$ images, resulting in sequences of L=16L=16k.
7.8 Empirical Comparison: Transformers vs. SSMs
We now provide an empirical comparison between attention-based Transformers and SSMs across a range of synthetic and real-world sequence modeling tasks. By evaluating performance on memory-intensive benchmarks like the Copy task, Induction Heads, Associative Recall, and LRA, the analysis highlights where each architecture excels or falls short, especially as sequence lengths grow or memory constraints tighten.
Copy task.   The synthetic Copy task [28] remains the most direct probe of a model’s raw memory capacity: the network is asked to reproduce an input string verbatim after a sentinel token. Recent results demonstrate a decisive advantage for attention-based Transformers over SSMs on both data-efficiency and length generalisation. Figure 20, revisited from [241], contrasts the strongest publicly reported results for Transformers and SSMs. Hard-ALiBi [241], a Transformer architecture exploiting positional encodings that bias query-key attention scores with a penalty proportional to their distance, reaches $\geq 95\%$ string-level accuracy on length-300 sequences after only $1\times 10^{4}$ training examples; conversely, Mamba-360M needs two orders of magnitude more (Figure 20–left). [^32] The authors use their standard 300-token setup to keep compute and optimisation schedules identical across models. When evaluated on out-of-distribution lengths (train on $\leq 50$, test on 300 / 1000), the Transformer preserves $\geq 90\%$ accuracy, while Mamba and other gated SSMs collapse to chance (Figure 20–right).
$10^{4}$$10^{5}$$10^{6}$0252550507575100100# Training examplesAccuracy (%)
[…]
Mamba-360MHard-ALiBi
Figure 20: Revisited from [241]. (Left) The authors train models to copy strings of length $\leq 300$ and evaluate string-level accuracy on strings of length $300$. Transformers train much faster than SSMs. Indeed, the Mamba-360M needs two orders of magnitude more w.r.t. Hard-ALiBi Transformer. (Right) The authors train models to copy on strings of length $\leq 50$ until all models are perfect in-distribution and evaluate string-level accuracy. Purple dashed line indicates maximum training string length and green dashed line indicates context window during training. Evaluating on longer inputs, the transformer models dramatically outperform the SSMs.
Induction Heads and Associative Recall.   Synthetic memory benchmarks such as Induction Heads (IH) and Associative Recall (AR) provide a minimal yet revealing playground for analysing how sequence models store and retrieve information. Because the training context is kept short while the evaluation context can reach millions of tokens, these tasks expose the limits of each architecture’s memory mechanism. We report in Table 6 the results in the original papers [151, 146]. All models are two‑layer variants trained on length-$256$ sequences for $2\times 10^{5}$ updates and evaluated on powers of two up to $L=1,048,576$ tokens. These tasks expose three qualitatively different memory regimes that map cleanly onto the architectures in Table 6. Plain linear SSMs such as S4D and GSS compress the entire past into a fixed-width latent state. As the number of distinct pairs approaches that width the representation collides, which explains the sharp drop to $\approx 35\%$ accuracy on IH even for 256–token contexts. H3 augments the recurrent path with a learned shift channel so that every incoming token is explicitly available in the next state’s slot, and combines this with multiplicative gating. The resulting mechanism behaves like a differentiable hash table that stores each token exactly once while using the gate for equality tests. This prior yields $99.8\%$ accuracy on AR and perfect IH, with the remaining $0.2\%$ error attributable to rare collisions when two identical keys land in the same step. Mamba replaces fixed convolutions with input–dependent transition matrices. By deciding at runtime which information to propagate, Mamba mimics the selective read–write behaviour of attention yet keeps both time and memory linear in $L$. Consequently it sustains $100\%$ accuracy from 64 up to 1,048,5761{,}048{,}576 tokens. When GPU memory allows, Transformers remain the simplest drop–in solution and excel at many-to-many interactions. H3 and especially Mamba become preferable for million-token contexts where linear memory is a hard requirement.
Table 6: Sequence‑level accuracy on Induction Heads (IH) and Associative Recall (AR)–results taken from [151, 146].
Model Parameters IH AR
Transformer [26] 0.10M 100% 100%
S4D [40] 0.10M 35.6% 86.0%
GSS [133] 0.10M 6.8% 78.0%
H3 [151] 0.15M 100% 99.8%
Mamba [146] 0.07M 100% 100%
Long Range Arena.   The LRA benchmark [35] comprises six tasks with sequence lengths ranging from 1,0241{,}024 to 16,38416{,}384 tokens, as described above. We compare the classical Transformer baseline against different SSMs architectures, reporting results from [42, 43] in Table 7.
Table 7: Exact test accuracy on the six Long-Range Arena tasks. Results are taken from table of the original papers [42, 43].
Task (length) Trans. Mega S5 LRU
ListOps (2 k) 36.37% 63.14% 62.15% 60.2%
Text (4 k) 64.27% 90.43% 89.31% 89.4%
Retrieval (4 k) 57.46% 91.25% 91.40% 89.9%
sCIFAR-10 (1 k) 42.44% 90.44% 88.00% 89.0%
Pathfinder (1 k) 71.40% 96.01% 95.33% 95.1%
Path-X (16 k) 7.00% 97.98% 98.58% 94.2%
Three main trends emerge from Table 7. Mega, which augments a standard self-attention block with a diagonal state-space kernel, delivers the top score on four of the six tasks—peaking at $91.25\%$ on the document–pair Retrieval task and $96.0\%$ on Pathfinder. The classical Transformer, by contrast, collapses to chance-level accuracy (7 %) on the 16k-token Path-X and lags by 20–50 pp everywhere else. S5 dominates the very long regime. When sequence length jumps to 16k, the continuous-time SSM of S5 reaches $98.58\%$ on Path-X. It also ties or edges Mega on the image and Pathfinder tasks while keeping $\mathcal{O}(L)$ memory and compute. LRU shows how far a linear RNN can go. With a purely recurrent core, stable-exponential parameterisation and a single $\gamma$-norm, LRU matches S5 within one percentage point on five tasks and even surpasses it on sCIFAR-10 (89.0 % versus 88.0 %). Taken together, these results suggest a clear trade-off. If memory footprint is no object, the Mega transformer still gives the best average accuracy. For workloads dominated by $\geq\!10\,000$-token sequences or tight latency budgets, however, S5 and the much simpler LRU offer comparable accuracy with strictly linear scaling and constant-time autoregressive generation.
In summary, Transformers outperform SSMs on data efficiency and generalization in the Copy task, requiring far fewer samples and maintaining high accuracy on longer sequences. In synthetic memory tests (IH, AR), advanced SSMs like H3 and Mamba match Transformer performance while scaling linearly with sequence length. On LRA benchmarks, Mega leads in accuracy, but SSMs like S5 and LRU offer competitive results with better efficiency, especially on very long sequences. Overall, Transformers are ideal when resources allow, while modern SSMs are better suited for long-context, resource-constrained settings.
@@
Benchmarks Transformers embracing Recurrence (Sec. 3) Deep State-Space models (Sec. 4) Novel RNNs architectures (Sec. 56)
Synthetic
Adding Problem [28] [235, 174] - [31, 168, 184, 190, 58]
Copy Task [28] [174, 206, 104, 123, 127] [39, 212, 44, 146] [168, 177, 206, 207, 223]
@@
Associative Recall [145, 123, 106] [151, 146, 74] [172, 200]
Computer Vision
sMNIST [30, 39, 40, 25, 42] [168, 177, 184, 189, 175, 190] [58, 234]
psMNIST [184, 186, 175, 190, 58, 223]
@@
Natural Language
WikiText103 [256] [84, 94, 85, 103, 114, 117, 123, 130, 99, 98, 102, 119, 143, 131] [25, 135, 154, 74] [168, 172, 206]
PTB[255] [117] [177, 190, 31, 162, 178, 58]
@@
Time Series
Eigenworms [289] [191, 186, 166]
@@ 8 Discussion and Future Directions
In this Section, we analyze some open problems and issues that have been recently pointed out in the works described in this survey, as well as highlighting possible future avenues for research. Despite the individual strengths of the architectures surveyed in previous sections, it is now increasingly evident that no single paradigm—be it pure Transformer, RNN, or SSM—can fully address the challenges posed by long-sequence modeling. Transformers offer excellent performance and global receptive fields, but suffer from quadratic complexity and require full sequence availability. RNNs and SSMs, on the other hand, offer efficient online processing and constant memory per time step, but are limited in terms of expressivity. These complementary strengths and weaknesses have naturally motivated the exploration of hybrid models that aim to combine the parallelism and global context handling of Transformers with the recurrence and efficiency of RNNs and SSMs. This section builds on this perspective, outlining both open issues and emerging design trends that point toward unified architectures capable of overcoming the fundamental trade-offs of current solutions.
8.1 Limitations of Transformers and Recurrent Models
The previous sections explored recent progress in Transformers and SSMs, highlighting the resurgence of recurrent mechanisms. Future research directions should be informed by an understanding of the current limitations of these model classes. There is potential for trade-offs and hybrid approaches, alongside a need to co-design models with the computational hardware they target.
Transformers.   A series of recent studies have evaluated Transformers across various tasks to uncover their limitations. Hanh et al. [303] demonstrated that Transformers struggle with recognizing simple patterns, such as parity or properly nested parentheses, when faced with long inputs. Peng et al. [304] further investigated these limitations, focusing on function composition. For instance, tasks involving multi-step reasoning over several facts remain challenging for Transformers. Although chain-of-thought prompting can alleviate some issues, it substantially increases the number of required tokens. Furthermore, Transformers have shown shortcomings in time series forecasting tasks, where simpler architectures have outperformed them [305]. Early theoretical work established that, under idealized assumptions such as infinite‑precision arithmetic and arbitrarily powerful feed‑forward subnets, transformers are “Turing complete” and can simulate a general‑purpose machine [306]. However, once we impose realistic constraints—most critically, limiting attention to hard (discrete) patterns—their power collapses into well‑studied Boolean circuit classes. Boolean circuits are acyclic networks of logical gates (e.g., AND, OR, NOT) whose depth and size determine how “parallel” or “sequential” a computation can be. Using circuit‐complexity techniques, Hahn [303] and Hao et al. [307] showed that transformers with hard attention correspond to non‑uniform $\mathrm{AC}^{0}$ circuits, which cannot even compute the majority function on $n$ bits. Allowing softer attention extends them to L‑uniform $\mathrm{TC}^{0}$ (the class of constant‑depth, polynomial‑size threshold circuits constructible in logarithmic space–informally, it can be thought of as the class of problems that can be solved with extremely parallel constant-depth computation), but Merrill & Sabharwal [308, 309] demonstrated that even this model cannot express inherently sequential computations. In particular, $\mathrm{NC}^{1}$‑hard state‑tracking problems—such as permutation composition—fall outside $\mathrm{TC}^{0}$ and thus lie beyond the reach of any realistic transformer model [310]. Permutation composition underpins fundamental tasks like playing chess, evaluating code, or tracking entities in a narrative; since these problems require carrying and updating an internal state step by step, transformers constrained by finite‑precision or limited attention cannot solve them in the worst case.
@@ 8.1 Limitations of Transformers and Recurrent Models
Recurrent models.   Despite the computational advantages of recurrence-based models, performances are not always on par with respect to vanilla attention when considering specific tasks, also in synthetic settings [151, 241]. Indeed, the capability of recurrent models to effectively solve a target task is strongly conditioned upon how efficiently the processed sequence is compressed into a finite-size state [146], which acts as an information bottleneck.
@@
The role of gating mechanisms turns out to be fundamental, given their influence on the model ability to control, in a context/input dependent manner, how information propagates or interacts along the temporal dimension. In fact, the relevance of an appropriate gating mechanism can be appreciated when considering that such a mechanism is present in many high-impact works described in this survey [237, 145, 171, 85, 94, 38, 75, 143], going well beyond the original case of LSTMs. Indeed, the Mamba model [146] shows how a selection mechanism (i.e., a term which refers to the mechanistic action of a model to select or ignore inputs and facilitate data interaction along the sequence length, inspired by the original gating in [28]) improves the capability of State-Space Models to solve tasks such as Selective copy or Induction Heads (see Section 7). In these task, content-aware reasoning is fundamental to be able to memorize relevant tokens and filter out the irrelevant ones. While the Linear Time Invariant (LTI) nature of previous SSMs (LSSL, S4 [40, 25]) allows to achieve efficient parallelizable solutions, their dynamics do not easily allow to select the correct information from the current context, making it harder to encode the information stored in the hidden state. Conversely, an appropriate selection mechanism introduces context-awareness, allowing the model to focus on or to filter out information. This element is also interesting in light of recent studies (see Section 5) suggesting that certain RNNs (equipped with linear recurrent layers interconnected by feedforward paths with multiplicative gating) might be implementing attention/selection “under the hood”, when trained to solve simple in-context learning tasks [171].
@@
However, despite such advantages brought by gating/selection mechanisms, the performance of recurrence-based solutions still falls short of fully closing the gap with attention-based models. Several recent works investigated such a recurrence expressivity gap [151, 241] with respect to attention. Indeed, inference in Transformers leverages the entire sequence processed so far, which, in autoregressive models, is stored in the KV-cache (see Section 3), to produce the final output. By definition, the memory usage of this approach increases over time, which becomes extremely expensive. At the cost of inefficiency, the attention mechanism possesses properties that standard stateful models are not able to embrace: explicitly recalling earlier tokens in the sequence (which influences their ability to copy from the input sequence, possibly by multiplying the attention scores with $V$) and comparing tokens across the sequence (which is achievable thanks to the quadratic attention matrix $QK^{\prime}$)). Jelassi et al. [241] exploit the copy task (see Section 7) to theoretically and empirically prove this property. In short, they show that from the theoretical point of view a simple Transformer model could be capable (when empowered by proper positional encodings and under some other mild assumptions) of copying strings of length that are exponential in the number of heads of the Transformer. This result relies on the ability of the Transformer to implement a mechanism of “storage” and retrieval of sequences of $n$ tokens ($n$-grams). Noticeably, the number of parameters depends only logarithmically on the input sequence length. The authors show that, conversely, any state-space model fails to solve the copy task, unless its latent state grows linearly with the sequence length. That is, SSMs cannot accurately copy strings with more bits than the size of the latent state.
@@
Such theoretical findings have been confirmed empirically. Indeed, Transformers are both much more efficient at learning to copy (indeed, they need $\approx 10^{4}$ training examples to learn the task, whereas Mamba [146] needs more than $\approx 10^{6}$ samples, as mentioned in Section 7.8) and also generalize better to longer inputs (when trained on strings $L=50$, they generalize to sequences with $L>1000$, whilst Mamba fails to generalize out of the trained context). This finding confirm the theoretical results, i.e., to be able to copy the entire input sequence, stateful models need to fully store it in its state space, which requires the memory to grow linearly with the sequence length. These results hold even when considering pre-trained models, both in copy tasks and retrieval benchmarks (e.g., Phonebook Lookup, see Section 7).